2. A uniform meter scale supported at the 40 cm mark is in equilibrium when a 50g mass is suspended at the 10 cm mark of the scale. Find the mass of the scale.
Answer #1

here,

mass suspended , m1 = 50 g

let the mass of scale be m2

equating the torques about the hinge

m1 * g * (40 - 10) - m2 * g * ( 50 - 40) = 0

50 * 30 = m2 * 10

m2 = 150 g

the mass of scale is 150 g